Question updated please need help in this questionGrassley Company completes these transactions during August of the current year (terms for all its credit sales are 2/10, n/3
Answer #1

1)

2)LO company payment : 2557 -557 ( returned goods ) = 2000 * 98% ( 2% discount ) = 1960